OpenStudy (anonymous):

let z=x+iy, showing details, find in terms of x and y: 1. Re[(z/barz), limz/z)

OpenStudy (anonymous):

\[z=x+iy,z ^{2}=\left( x+iotay \right)^{2}=x ^{2}+\iota ^{2}y ^{2}-\iota2xy=x ^{2}-y ^{2}-\iota2xy\] \[\left( 1+\iota \right)^{2}==1+\iota ^{2}+2\iota=1-1+2\iota=2\iota \] \[\left( 1+\iota \right)^{4}=\left( 2\iota \right)^{2}=4\iota ^{2}=-4\] \[\left( 1+\iota \right)^{16}=\left( -4 \right)^{4}=256\] (1+i)^16 z^2=256[(x^2-y^2)-i2xy] \[\mathbb{R} \left[ \left( 1+i \right)^{16}z ^{2} \right]=256\left( x ^{2} -y ^{2}\right)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but how did -i2xy dissapear

OpenStudy (anonymous):

please

OpenStudy (anonymous):

2. Re(z/barz), lim(z/z) barz is conjugate of z.